Find the coorddinates of the point of intersection of the straight lines 2x-3y=1 and 5y-x=3 and determine the angle between them

Grade:10

3 Answers

Aditya Gupta
4 years ago
to find the point of intersection simply solve the two eqns together. 2x-3y=1 and 5y-x=3 gives x= 2 and y= 1.
so coordinates of the point of intersection= (2, 1)
now, the formula for angle between 2 lines is 
theta= arctan[|(m1 – m2)/(1+m1*m2)|], where m1 and m2 are the slopes of the lines.
here, m1= 2/3 for 2x-3y=1 and m2= 1/5 for 5y-x=3.
so, theta= arctan[|(m1 – m2)/(1+m1*m2)|]= arctan(7/17) which is approx 22.38 degrees.
kindly approve :)
Vikas TU
4 years ago
Dear student 
in equation 1 , a1 = 2 , b1 = -3 , 
in equation 2 a2 = 5 , b2 = -1 
a1/a2 is not equal to b1 / b2 
so, this is the condition of line to intersect , 
and angle between them = 22.38 degree.
Pawan Kumar Karela
4 years ago
Given equation    2x-3y=1   (1)
                            5y-x=3     (2)
we use the elimination method to get the intersection point.
therefore multiply equation (2) by 2 and then add it to equation (1) to eliminate variable y.
                          2x-3y=1     (3)
                         -2x+10y=6  (4)
adding (3) and (4) we get 
                            7y=7
                           y=1      (*)
              use value of y to get x
                     2x-3*1=1
                        2x=1+3
                         x=2
         therefore point of intersection is (2,1)
 
to find angle between these two lines we use formula
                                  tan\Theta =\left | \frac{m1-m2}{1+m1*m2} \right |
from equation (1)  2x-3y=1...... implies y=(2/3)x-1
                                                      therefore m1=2/3
similarly for equation (2)  y=1/5+3/5
                                       therefore m2=1/5
               hence tan\Theta =\left | \frac{(2/3)-(1/5)}{1+(2/3)*(1/5)} \right |
                           therefore on solving we get
                                      tan\Theta =\left | \frac{7}{17}\right |
                                       \Theta =tan^-1\left | \frac{7}{17}\right |